Create a proposal to start a new company. You will sell one product only and will need to have a variety of fixed and variable costs in the make-shift company. Fictionalize a data set and use this data to generate a CVP analysis showing both the break-even point and also the number of units to be sold in order to earn a $10,000 profit.
Answer #1
Suppos there is product A whose selling price is $ 100
its variable Cost is $60 and Fixed Cost is $10000
Selling Price 100
Variable Cost 60
Contribution 40
Fixed Cost 10000
Its contribution Margin is 40%
Break Even Point in unit = Fixed Cost / Contribution per unit
=10000/40
250
BEP in $ =250*100
25000
No of units required to be sold to earn $15000 profit
=(10000+15000)/40
625
$ sales required to be sold to earn $15000 profit =625*100
62500